in a triangle ABC,if angle A+angle B =150 degree and angle B+angle C=100.find the measure of each angle of the triangle
Answer by Cromlix(4381)

If angle A + angle B = 150 degrees
then angle C = 180 - 150 = 30 degrees.
If angle B + angle C = 100 degrees
then angle A = 180 - 100 = 80 degrees.
Then angle B = 180 - (30 + 80)
= 180 - 110
= 70 degrees
Angle A = 80 degrees
Angle B = 70 degrees
Angle C = 30 degrees
Hope this helps.
:-)